the book only says find the solution and i can't find an example to work from
help
x^4-5x^2+4=0

Found 2 solutions by Alan3354, rapaljer:
Answer by Alan3354(30993)

It's a quadratic in x^2
Sub y for X^2:
y^2 - 5y + 4 = 0
Can you finish it?
email via the Thank You note for more help.

Answer by rapaljer(4667)

x^4-5x^2+4=0
(x^2-4)(x^2-1)=0
(x-2)(x+2)(x-1)(x+1)=0

There are 4 solutions:
x-2=0
z=2

x+2=0
x=-2

x-1=0
x=1

x+1=0
x=-1

To see my explanation of this topic on my own website, please go to "www2.seminolestate.edu/rrapalje". Look for the link on my Homepage called "Basic, Intermediate, and College Algebra: One Step at a Time." Look under "Basic Algebra" for Chapter 2, Solving Quadratic Equations.

I think you will find the explanations easier to understand than traditional textbooks. My own students did!!

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus